OpenStudy (anonymous):

secx^2-1/sinx=sinx/1-sin^2x

OpenStudy (mathmale):

aatka: Hope you won't take my input too negatively: I strongly suggest that you include with all these equations and symbols the directions we're supposed to follow in "solving" this problem. Secondly, I'd suggest that when dividing by expressions such as 1-(sin x)^2, you enclose those expressions in parentheses so that there's no doubt as to what your divisor includes and does not include: [1-(sin x)^2]. Please type in the instructions for this problem; perhaps then I could give you some help.

OpenStudy (anonymous):

Is this your problem \[ \frac{\sec ^2(x)-1}{\sin (x)}=\frac{\sin (x)}{1-\sin ^2(x)} \] To show that they are equal

OpenStudy (anonymous):

Notice that \[ \frac{\sec ^2(x)-1}{\sin (x)}=\tan (x) \sec (x) \] Can you show that the second side is also equal to \[ \tan (x) \sec (x) \]

OpenStudy (anonymous):

You can show easily that \[ \frac{\sin (x)}{1-\sin ^2(x)}=\tan (x) \sec (x) \]